OpenStudy (anonymous):

When the coffee is brewed according to directions, a pound of coffee beans yields 50 cups of coffee (4 cups = 1 qt). How many kg of coffee are required to produce 100 cups of coffee?

OpenStudy (anonymous):

\[\frac{ cupsofcoffee }{ 1 }\times \frac{ pound of coffee }{ cups ofcoffee }\times \frac{ kg of coffee }{ poundofcoffee}\] Can you figure out without numbers? Try to start with given values that are usually in the last sentence.